# Discretion in the Loop: Human Expertise in Algorithm-Assisted College Advising

### v1

@@ 14 Unregistered - - - - - 12 12
Table 3: Counts of specified reason categories for each intervention type mentioned during advising meetings, indicating how frequently each reason was associated with a particular intervention, for Intervention 2 (Correct Schedule), Intervention 4 (Progression Issues), Intervention 5 (Adjust Course Balance), Intervention 9 (Address Financial Issues), Intervention 10 (Address Non-Financial Collegiate Requirements), and Intervention 16 (Schedule Appointment).
7 Heterogeneity in Student Outcomes
Beyond studying expertly targeted actions and its mechanisms, in this section, we explore the broader question of impact: In what ways does discretionary decision making affect student outcomes and equity? Motivated by recent work on the equity implications of algorithmic interventions in education [23, 24], we focus our analysis on how treatment effects (on graduation outcomes) vary across subgroups defined by (1) the primary advisor assigned to students; and (2) student demographics, specifically race and ethnicity. To estimate these heterogeneous treatment effects, we fit logistic regression models, as well as causal forest models to account for potential non-linear interactions. All models are trained on data from 745 students (392 in the treatment group and 353 in control). [^27] For these regressions, we exclude students identified as multiracial and those categorized as non-resident aliens, as these groups are substantially smaller in size compared to other demographic categories (see Appendix A.6.1 for a complete list of exclusions).
Across advisors, we observe some heterogeneity in graduation outcome effects (see Table 18), though the differences are not statistically significant. However, a qualitative analysis of advisor comments in §7.1.2 suggests variations in advising approaches. Synthesizing both analyses, we find preliminary evidence that student outcome variations may be associated with advising style: students assigned to advisors with a holistic approach tended to fare better than those with academically-focused advisors.
The original MAAPS study (§3.1.1) noted potentially greater treatment effects for Black students. [^28] After six academic years, Black students experienced a more positive treatment effect on graduation outcomes from the MAAPS intervention than non-Black students [65]. Building on this, we conduct a more systematic analysis of heterogeneous treatment effects across all race and ethnicity groups. We fail to find significant heterogeneous effects by race-ethnicity after correcting for multiple hypothesis testing (see results in Table 20). We remark, however, that the original MAAPS study is not well powered to detect small differences in treatment effect (see discussion in Appendix A.6.5), and the lack of statistical significance need not imply the absence of heterogeneous treatment effects.
7.1 Heterogeneous Effects Across Advisors
We look for heterogeneous effects across advisors.
7.1.1 Quantitative Analysis
Each student in the MAAPS study treatment group was assigned to one of the five different advisors. For each of the advisors $j\in\{1,\ldots,5\}$, let $A_{ij}$ denote whether student $i$ was assigned to advisor $j$ [^29] Some students had interactions with multiple MAAPS advisors. For these students, we let $A_{ij}=1$ if most of student $i$’s interactions are with advisor $j$ and we set $A_{ij^{\prime}}=0$ for all $j^{\prime}\neq j$., $Z_{i}\in\{0,1\}$ denote whether the student has successfully graduated, and $T_{i}\in\{0,1\}$ indicate whether the student was assigned to the treatment group. We fit the following logistic regression to model the graduation outcome, [^30] We additionally considered exploratory analyses for the cumulative GPA outcomes. However, due to substantial selection bias in those models (only enrolled students have a GPA), regression results may not provide reliable or generalizable inferences. controlling for the baseline student attributes (gender, high school performance reflected in the composite ACT, financial status in terms of expected financial contribution, and academic preparation measured by other credit hours, which we describe further in Appendix A.1.1):
$\begin{split}\text{logit}&\big{(}\mathbb{P}(Z_{i}=1)\big{)}=\,\beta_{0}+\beta_{1}T_{i}+\sum_{j}\theta_{j}A_{ij}+\gamma_{1}\text{(Gender)}_{i}\\ &+\gamma_{2}\text{(ACT)}_{i}+\gamma_{3}\text{(EFC)}_{i}+\gamma_{4}\text{(Hours\_Other)}_{i}\end{split}$ (1)
We do not find significantly non-zero heterogeneous treatment effects across advisors (see Appendix Table 18). However, both logistic regression and causal forest models produce a consistent ranking of advisor-specific effects (Appendix Table 19). In the logistic regression model, Advisor 3 has the highest advisor-specific coefficient $\theta_{3}=0.51$, followed by Advisors 2 and 1 ($\theta_{3}$ of 0.02 and $-0.25$ respectively). Similarly, the causal forest analysis estimates the highest heterogeneous treatment effect for Advisor 3, with an effect size of 0.11. The ensuing qualitative analysis sheds light on potential connections to differences in advising style.
7.1.2 Qualitative Analysis
Figure 2: Normalized distribution of qualitative codes (meeting topics) for each MAAPS advisor. Advisors 1 and 2 focus on academic topics, while Advisor 3 emphasizes relationship-building and nonacademic discussions during student meetings. The Generic Suggestion code here refers to the advisor asking the student to follow up with them later without specifying a reason (see full details in Appendix A.4.1).
Despite the lack of significant effects in the quantitative analysis, the qualitative analysis of the advisors’ comments (§6.1.1) reveals notable differences in their approaches to advising. In particular, we highlight two distinct advising styles that primarily differ in the scope of topics discussed in advisor-student interactions. [^31] We note that our analysis is limited to advisors’ written comments, which might not reflect the complete range of applied advising strategies. The first style (reflected in the comment of Advisors 1 and 2) is academic-oriented, in which the advisor adheres to a focused set of actions and topics in meetings including academic performance, degree progression and requirements. On the other hand, the second style (Advisor 3) is holistic, with broader engagement in non-academic matters and student’s interests and emotional states. [^32] Advisors 4 and 5 are excluded from our analysis due to the limited number of comments, which did not provide enough data for a meaningful comparison. Figure 2 shows the full distribution of subject matter codes for each MAAPS advisor.
To illustrate the academic-focused style of advising, we turn to Advisor 1, whose comments consistently center on academic planning, frequently including references to degree audits, major maps, and specific course planning. Advisor 1’s comments follow a structured pattern across most student meetings, reflecting a more routine approach to academic advising: [^33] Advisor 1 utilized a standardized comment template in 76% of all meetings (647 out of 849 total comments). Of these repeated comments, 440 include ”Additional comments”, although only 44 specifically addressed unique student concerns.
"Student came to the office to discuss schedule for SP’18 semester, printed and reviewed degree audit, student was shown major map, reviewed EAB for success markers, and notified student…"
Interestingly, we notice that Advisor 1 is the only advisor who explicitly refers to the use of algorithmic tools (“reviewed EAB for success markers”) during advisement.
When Advisors 1 and 2 do personalize their advice to the student, their scope of consideration is also largely constrained to the student’s academic status. For example, Advisor 2 writes:
"Student wants to change major from bio to exercise science… I advised that they could possibly do mini-mester for spring in order to compensate for credits lost towards degree."
Advisor 3 exemplifies a more holistic advising style. The comments left by Advisor 3 reflect a broader engagement with the student as a whole person, addressing not only academic planning, but also emotional well-being, personal motivations, and long-term goals. The advisor is also more likely to note and adjust their advice to the specific concerns and needs that students bring up during each meeting, even if these lie outside the scope of academics.
In one example, Advisor 3 encourages a student to reflect more deeply on their desire to help others, guiding them to explore opportunities that align with their interests:
"The student wants to help people, but I challenged them to think about why and how. I advised them to look into some organization fairs where they can start getting experience working with people."
Advisor 3’s comments also reveal a pattern of empathic support and thoughtful challenge, which is also reflected in a significantly higher proportion of relationship-building qualitative codes applied to each student meeting (see Figure 2). For example, Advisor 3 responds to a student experiencing anxiety about their future by acknowledging their concerns, exploring possible academic paths, and pointing them toward mental health resources:
“Student is very anxious about GSU, college and career in general… I challenged them to get exposure to the field. They did ask about the counseling center, for which I provided them with resources.”
The structure and content of these conversations vary considerably by advisor. As shown in Table 4, Advisors 1 and 2 have relatively high average Jaccard Similarity Index (JSI) scores,343434 For a pair of meetings, the Jaccard Similarity Index (JSI) is defined as the ratio of the intersection to the union of the codes applied to each meeting. For each advisor, we calculate the average JSI for all pairs of their meetings to assess how similar their comments were overall, with a higher average indicating less diversity in comments. reflecting a repetitive use of qualitative codes across sessions, consistent with a structured advising approach. In contrast, Advisor 3 has a notably lower JSI value, indicating diversity in the topics and strategies employed, which aligns with a personalized advising style.
Advisor # Comments Average JSI
Advisor 1 843 0.3006
Advisor 2 364 0.3791
Advisor 3 395 0.1914
Table 4: Average pairwise Jaccard Similarity Index (JSI) on qualitative codes applied in the meetings for each advisor.
The quantitative analysis of heterogenous treatment effects suggests that Advisor 3 has a more positive impact on student outcomes than Adviors 1 and 2. This observation may reflect the advantages of a more holistic advising style compared to the academic-oriented approach for this particular college advising context. In our setting, it is also important to note the limited power to detect heterogeneous effects and the potential for unmeasured effects, such as student satisfaction or self-efficacy, which may still be influenced by the advising styles but are not captured in the graduation-based quantitative analysis. Future work is required to further investigate whether certain advising styles lead to more student success.
7.2 Heterogeneous Effects by Student Demographics
To examine heterogeneity in treatment effects across student subgroups, we estimate a logistic regression model predicting the binary graduation outcome $Z_{i}\in\{0,1\}$, where $Z_{i}=1$ indicates that student $i$ graduated. The main treatment indicator $T_{i}\in\{0,1\}$ denotes assignment to the treatment group. For the independent variables, we include student race-ethnicity indicators (Black, Asian, Hispanic [^35] The indicator variable White was omitted from the model to avoid multicollinearity.), their interactions with treatment, and baseline covariates (as controls), using the same specification as Model (1).
$\begin{split}\text{logit}\big{(}\mathbb{P}(Z_{i}=1)\big{)}=\,&\beta_{0}+\beta_{1}T_{i}+\sum_{r\in\mathcal{R}}\left(\theta_{r}\cdot\text{Race}_{r,i}+\delta_{r}\cdot T_{i}\cdot\text{Race}_{r,i}\right)\\ \quad+\gamma_{1}\text{(Gender)}_{i}+&\gamma_{2}\text{(ACT)}_{i}+\gamma_{3}\text{(EFC)}_{i}+\gamma_{4}\text{(Hours\_Other)}_{i}+\epsilon_{i}\end{split}$ (2)
We focus specifically on subgroup effects by race, motivated by Georgia State University’s institutional commitment to improving educational access and equity for historically underrepresented students [61], as well as the broader concerns that early alert and predictive analytics systems can unintentionally hurt educational outcomes for minority students [23]. We note that race is not used as an input in the predictive analytics system deployed at GSU.
We do not find statistically significant interaction terms between treatment and race variables, after correcting for multiple hypothesis testing (see results in Table 20). This may not be an indication of the lack of heterogeneous treatment effects, as the MAAPS experiment may not be well-powered to detect small heterogeneous treatment effects. We run a power analysis in Appendix A.6.5 and find that, assuming a desired power of 80% and a sample size of 745 students from our experiments, the minimal additive detectable effect is 18%, meaning that our study on the MAAPS dataset can only detect a large increase in the graduation rate. Furthermore, from Table 17, we can estimate that in order to achieve significant increase of the average graduation of $\sim$ 5% within the subgroup of Black students, we would require a total population sample size of approximately 10,000 students.
Additionally, to assess whether advisors’ intervention choices were systematically influenced by students’ race and ethnicity, we conduct a series of logistic regression analyses. For each of the five significant interventions (listed in §5.2.1), we estimate two models (see Appendix A.6.8): the first that included race-ethnicity variable alongside all other baseline student characteristics (listed in Appendix A.1.1), and the second model that excludes this variable. Across all cases, we find no difference in predictive performance between these two classifiers (see Table 22), suggesting that, conditional on observed student characteristics, advisor decisions to select particular interventions during meetings were not meaningfully shaped by student race and ethnicity.
Together, these findings highlight the need for further investigation into how discretionary decision making interacts with equity in student support.
@@ 8.3 Functional and Normative Roles of Human Discretion in Institutional Decision Making
On the negative side, discretion can hinder decision making, particularly in settings where decision quality is thought to depend mostly on prediction accuracy [41] (though this is not our setting). In some predictive contexts, algorithms may outperform human forecasters [20] and so lead to improvements in decision quality. Even when human decision-makers have access to information beyond what the algorithm does [5], studies show that people often misjudge when to override or defer to algorithmic advice, leading to over-reliance or under-reliance [9]. Moreover, discretion can introduce bias, as demonstrated by the case of a pretrial risk assessment tool in Kentucky, where judges were more likely to override low-risk scores for Black defendants than for white defendants [3]. Our finding that the MAAPS study is underpowered to detect heterogeneous effects for minority students is thus a notable limitation, as it constrains our ability to assess whether discretion in advising contributed to student outcome disparities in unintended ways.
@@ Acknowledgments.
We would like to thank the hard work of the entire MAAPS project team at GSU. The efforts of the MAAPS advisors, particularly the diligence and dedication of Emily Buis, were instrumental in developing a rich source of data that made this work possible. We thank Timothy Renick for his visionary leadership at NISS and for sharing insights that provided valuable context and inspiration for this research collaboration. We also thank Tim Fulton, a prior staff member at NISS, for his support, as well as Andy Zhang for his research assistantship in the initial stages of this study. We thank Manish Raghavan for valuable discussions on his earlier work.
@@ A.2 Theoretical Framework
Next, we show both of these models meet our assumption that SCMs are causally minimal. Causal minimality is equivalent to that, for all causal parents of $Y$, $X\in\mathbf{PA}_{Y}$, $X\not\perp Y\mid\mathbf{PA}_{Y}\backslash\{X\}$ [57, p.109]. Accordingly, for $\mathcal{M}_{1}$, we show that both $(1)Y_{t+1}\not\perp U_{t}|A_{t}$ and $(2)Y_{t+1}\not\perp A_{t}|U_{t}$. All probabilities are calculated by summing over the relevant rows of Table 5.
@@
    Parents do/see is that conditioning and intervening on parent values of a variable are interchangeable [11, p.32]. It is true for all DAGs generated by SCMs [11, p.33].
@@
Recall that causal minimality is equivalent to that, for all causal parents of $Y$, $X\in\mathbf{PA}_{Y}$, $X\not\perp Y\mid\mathbf{PA}_{Y}\backslash\{X\}$ [57, p.109].
@@
∗∗∗ Recall parents do/see is that conditioning and intervening on parent values of a variable are interchangeable [11, p.32]. It is true for all DAGs generated by SCMs [11, p.33].
@@ 18. See primary program advisor 49 6 7 1 3
19. Academic map/degree planner 635 7 19 14 6
20. Other (Describe in comments) 1810 41 70 10 23
@@ A.3.5 Full Results
Additionally, the top 5 most significant interventions at GPA binarization thresholds ranging from 2.5 to 4.0 are included in Table 13. As the table demonstrates, Interventions 4 (Progression Issues) and 16 (Schedule an Appointment) are significant for all tests, while Intervention 2 (Correct Schedule) and Intervention 20 (Other) are always in the top 5 except for in predicting a perfect GPA of 4.0 (which would be expected due to the especially high bar in addition to the reduced power of the test when most people’s outcomes do not differ).
@@ A.3.6 Correlations Between Interventions
For persistence outcomes, only one intervention (Intervention 17, Recommend Schedule) is significant, so no correlations warrant further attention. For GPA outcomes, Intervention 4 (progression issues) shows moderate correlation with each of the other significant interventions: Intervention 2 (Correct Schedule, $r=0.44$, $p=6.4\times 10^{-117}$), Intervention 16 (Schedule Appointment, $r=0.35$, $p=1.31\times 10^{-72}$), and Intervention 20 (Other, $r=0.30$, $p=1.5\times 10^{-53}$), while the remaining pairs of interventions show only weak correlation. Notably, Intervention 20 (Other) predicts a GPA $\geq 3.5$ positively while Intervention 4 (Progression Issues) predicts it negatively, making it unlikely that their conditional dependence with outcomes reflects joint causal effects or targeting. Meanwhile, qualitative analysis supports the interpretation that Interventions 2 (Correct Schedule) and 16 (Schedule Appointment) are informed by different types of non-algorithmic information than Intervention 4 (Progression Issues) (Table 3), suggesting they are separately targeted. However, given the moderate correlations, it remains possible that Interventions 2 (Correct Schedule) and 4 (Progression Issues), as well as Interventions 4 (Progression Issues) and 16 (Schedule Appointment), are either jointly non-algorithmically targeted or jointly influence GPA outcomes.
@@ A.4 Exploratory Qualitative Analysis
In our analysis, codes served as succinct, essence-capturing labels that represent meaningful patterns in the data [67]. We organized codes hierarchically, with high-level categories capturing key differences in advisor style and aims (e.g. discussing a topic open-endedly, making directive suggestions to the student, or building rapport with the student) and finer-grained codes differentiating interactions within these categories (for instance, the discussion topic). We report the frequency of these applied codes by each advisor in Table 14. Our complete codebook with representative quotes is provided below.
@@
    • Scheduling - discussing course schedules, registration timelines, or planning for upcoming semesters:
      “Student wants to change their schedule for the spring and take some courses online. We discussed several available options."
    • Academic administrative - addressing administrative issues like degree requirements, deadlines, navigating transfers, or institutional processes:
      “Reminded a student about the deadline to submit AP, DE, and/or IB transcripts."
    • Major/minor selection - discussing the student choice of major/minor:
      “Student is thinking about changing majors to finance. They like numbers, but also enjoy being creative. I gave them the Focus 2 Assessment to help identify their strengths."
    • Academic performance - reviewing grades, progress, or challenges in coursework:
      “Student wanted to withdraw from [ECON XXX], as their first exam was a 40%; they admitted they just didn’t study. We talked through their degree requirements and they understand the importance of taking this class seriously now."
    • EAB/algorithmic review - reviewing algorithmically-suggested or institutional benchmarks for the student, like EAB’s Success Markers or Academic Improvement Plan warnings:
      “I have printed and reviewed degree audit, reviewed EAB for success markers, and notified student.”
@@
    • Career - talking about career aspirations, opportunities, career fairs, or preparation:
      “Student is interested in going to law school. I explained popular majors for pre-law route such as POLS, IDS, and Philosophy.”
    • Financial aid - discussing applying for or maintaining financial aid packages/status:
      “Student mentioned they will go to financial aid due to their refund not going to the right bank. This is vital as they have not been able to get their school books yet."
    • Financial trouble - discussing problems students are having with finances:
      “Student is on the drop-list for non payment. I called the student to warn them and to come up with a plan, and talk about reinstatement."
    • Other nonacademic - discussing any other nonacademic matter, such as extracurriculars, housing/commuting, or family, medical concerns:
      “Student is living on campus, has joined several social organizations, and we talked about work-life balance."
@@
    • Follow up with advisor - the advisor asks students to check in with them, schedule a meeting with them, or follow up in the future with updates about a specifically academic matter:
      “Asked student to sit down together and discuss each hold, and help them figure out where to go to get them resolved."
    • Follow up with another person - the advisor asks students to speak with another person about a specifically academic matter:
      “I encouraged them to talk to their instructors of [philosophy course] and [film course], to ask about any tips for the remaining of the semester."
    • Guiding/steering the student - the advisor tells a student how to alter their program of study, for instance by changing major, recommending a schedule, or adding/dropping a class:
      “I told them I’m nervous about them having GEOG and MATH together and suggested to revise the schedule; however, they were persistent about completing them both in the Spring semester."
@@
    • Follow up with advisor - the advisor asks students to check in with them later on or lets them know that they can check in with future questions about specifically nonacademic matters:
      “Student will return in the near future with top 5 schools they are considering for medical school."
    • Other suggestion - the advisor makes any other sort of suggestion about specifically non-academic matters, for instance referring students to a campus office or recommending a student pursue an extracurricular:
      “Advisor confirmed that student can still fulfill law school requirements with courses and completed what-if analysis. Advised student to visit career services for further confirmation and to complete assessments. "
@@
    • Follow up with advisor - the advisor asks students to check in with them later / follow up later, without mentioning any specific topic:
      “Student will continue to explore their options as far as choosing a major. Asked the student to follow up if they have any additional questions or concerns."
@@
    • Empathy/sympathy - expressing concern, understanding, or emotional support for the student’s circumstances or challenges:
      “I am very proud of you for completing your 1st year of college and looking forward to your success in the upcoming academic year."
    • Student affect - observing and noting the student’s emotional state or behavior during the interaction:
      “Student really likes GSU, joined Greek life. I asked about their school and social life balance was going, they said good."
@@
    • Changing the student’s schedule - directly altering a student’s schedule on their behalf:
      “I looked through student’s degree planner and their schedule, and adjusted it due to needing [MATH XXX]."
    • Contacting student’s parents - communicating directly with a student’s family members to address concerns or provide important information:
      “Spoke with student’s parent explaining to them that their kid needs to complete their FAFSA for the 17/18 academic year."
    • Speaking with department/instructor/administrator - The advisor reaches out to a student’s department, instructor, or an administrator on behalf of them.
      “Student had an issue with one of the professors. I walked them over to the Ombudsperson office to meet with Assistant Ombudsperson […] to discuss the issue."
    • Signing / filling out form for student - The advisor signs or completes a form on behalf of the student:
      “Student is within transient guidelines as acknowledged on the document. I completed the form, signed, and sent it to the registrar’s office for further processing."
@@
Table 14: Frequencies of the applied codes by each advisor during student-advising meetings, based on our exploratory qualitative analysis. Notable differences in code distributions across advisors highlight distinct advising styles.

### v2

@@ 14 Unregistered - - - - - 12 - 12
Table 3: Counts of specified reason categories for each intervention type mentioned during advising meetings, indicating how frequently each reason was associated with a particular intervention, for Intervention 2 (Correct Schedule), Intervention 4 (Progression Issues), Intervention 5 (Adjust Course Balance), Intervention 9 (Address Financial Issues), Intervention 10 (Address Non-Financial Collegiate Requirements), Intervention 16 (Schedule Appointment), and Intervention 20 (Develop Degree Planner). Results for Intervention 17 (Recommend Schedule) are presented separately in Appendix A.5.3 since we look for signs of non-algorithmic information in schedule customizations rather than reasons to recommend a schedule.
@@ 6.3 Functional and Normative Roles of Human Discretion in Institutional Decision Making
On the negative side, discretion can hinder decision making, particularly in settings where decision quality is thought to depend mostly on prediction accuracy [38] (though this is not our setting). In some predictive contexts, algorithms may outperform human forecasters [18] and so lead to improvements in decision quality. Even when human decision-makers have access to information beyond what the algorithm does [4], studies show that people often misjudge when to override or defer to algorithmic advice, leading to over-reliance or under-reliance [7]. Moreover, discretion can introduce bias, as demonstrated by the case of a pretrial risk assessment tool in Kentucky, where judges were more likely to override low-risk scores for Black defendants than for white defendants [2].
@@ Acknowledgments.
We would like to thank the hard work of the entire MAAPS project team at GSU. The efforts of the MAAPS advisors, particularly the diligence and dedication of Emily Buis, were instrumental in developing a rich source of data that made this work possible. We thank Timothy Renick for his visionary leadership at NISS and for sharing insights that provided valuable context and inspiration for this research collaboration. We also thank Tim Fulton, a prior staff member at NISS, for his support, as well as Andy Zhang for research assistantship in the preliminary stage of this study. We thank Manish Raghavan for valuable discussions on his earlier work.
@@ A.2 Theoretical Framework
Next, we show both of these models meet our assumption that SCMs are causally minimal. Causal minimality is equivalent to that, for all causal parents of $Y$, $X\in\mathbf{PA}_{Y}$, $X\not\perp Y\mid\mathbf{PA}_{Y}\backslash\{X\}$ [53, p.109]. Accordingly, for $\mathcal{M}_{1}$, we show that both $(1)Y_{t+1}\not\perp U_{t}|A_{t}$ and $(2)Y_{t+1}\not\perp A_{t}|U_{t}$. All probabilities are calculated by summing over the relevant rows of Table 4.
@@
    Parents do/see is that conditioning and intervening on parent values of a variable are interchangeable [9, p.32]. It is true for all DAGs generated by SCMs [9, p.33].
@@
Recall that causal minimality is equivalent to that, for all causal parents of $Y$, $X\in\mathbf{PA}_{Y}$, $X\not\perp Y\mid\mathbf{PA}_{Y}\backslash\{X\}$ [53, p.109].
@@
∗∗∗ Recall parents do/see is that conditioning and intervening on parent values of a variable are interchangeable [9, p.32]. It is true for all DAGs generated by SCMs [9, p.33].
@@ 18. See primary program advisor 49 6 7 1 3
19. Other (Describe in comments) 635 7 19 14 6
20. Develop degree planner 1810 41 70 10 23
@@ A.3.5 Full Results
Additionally, the top 5 most significant interventions at GPA binarization thresholds ranging from 2.5 to 4.0 are included in Table 12. As the table demonstrates, Interventions 4 (Progression Issues) and 16 (Schedule an Appointment) are significant for all tests, while Intervention 2 (Correct Schedule) and Intervention 20 (Develop Degree Planner) are always in the top 5 except for in predicting a perfect GPA of 4.0 (which would be expected due to the especially high bar in addition to the reduced power of the test when most people’s outcomes do not differ).
@@ A.3.6 Correlations Between Interventions
For persistence outcomes, only one intervention (Intervention 17, Recommend Schedule) is significant, so no correlations warrant further attention. For GPA outcomes, Intervention 4 (progression issues) shows moderate correlation with each of the other significant interventions: Intervention 2 (Correct Schedule, $r=0.44$, $p=6.4\times 10^{-117}$), Intervention 16 (Schedule Appointment, $r=0.35$, $p=1.31\times 10^{-72}$), and Intervention 20 (Degree Map, $r=0.30$, $p=1.5\times 10^{-53}$), while the remaining pairs of interventions show only weak correlation. Notably, Intervention 20 (Degree Map) predicts a GPA $\geq 3.5$ positively while Intervention 4 (Progression Issues) predicts it negatively, making it unlikely that their conditional dependence with outcomes reflects joint causal effects or targeting. Meanwhile, qualitative analysis supports the interpretation that Interventions 2 (Correct Schedule) and 16 (Schedule Appointment) are informed by different types of non-algorithmic information than Intervention 4 (Progression Issues) (Table 3), suggesting they are separately targeted. However, given the moderate correlations, it remains possible that Interventions 2 (Correct Schedule) and 4 (Progression Issues), as well as Interventions 4 (Progression Issues) and 16 (Schedule Appointment), are either jointly non-algorithmically targeted or jointly influence GPA outcomes.
@@ A.4 Exploratory Qualitative Analysis
In our analysis, codes served as succinct, essence-capturing labels that represent meaningful patterns in the data [61]. We organized codes hierarchically, with high-level categories capturing key differences in advisor style and aims (e.g. discussing a topic open-endedly, making directive suggestions to the student, or building rapport with the student) and finer-grained codes differentiating interactions within these categories (for instance, the discussion topic). We report preliminary frequencies of these applied codes by each advisor in Table 13. Our complete codebook is provided below.
@@
    • Scheduling - discussing course schedules, registration timelines, or planning for upcoming semesters
    • Academic administrative - addressing administrative issues like degree requirements, deadlines, navigating transfers, or institutional processes
    • Major/minor selection - discussing the student choice of major/minor
    • Academic performance - reviewing grades, progress, or challenges in coursework
    • EAB/algorithmic review - reviewing algorithmically-suggested or institutional benchmarks for the student, like EAB’s Success Markers or Academic Improvement Plan warnings
@@
    • Career - talking about career aspirations, opportunities, career fairs, or preparation
    • Financial aid - discussing applying for or maintaining financial aid packages/status
    • Financial trouble - discussing problems students are having with finances
    • Other nonacademic - discussing any other nonacademic matter, such as extracurriculars, housing/commuting, or family, medical concerns
@@
    • Follow up with advisor - the advisor asks students to check in with them, schedule a meeting with them, or follow up in the future with updates about a specifically academic matter
    • Follow up with another person - the advisor asks students to speak with another person about a specifically academic matter
    • Guiding/steering the student - the advisor tells a student how to alter their program of study, for instance by changing major, recommending a schedule, or adding/dropping a class
@@
    • Follow up with advisor - the advisor asks students to check in with them later on or lets them know that they can check in with future questions about specifically nonacademic matters
    • Other suggestion - the advisor makes any other sort of suggestion about specifically non-academic matters, for instance referring students to a campus office or recommending a student pursue an extracurricular
@@
    • Follow up with advisor - the advisor asks students to check in with them later / follow up later, without mentioning any specific topic
@@
    • Empathy/sympathy - expressing concern, understanding, or emotional support for the student’s circumstances or challenges
    • Student affect - observing and noting the student’s emotional state or behavior during the interaction
@@
    • Changing the student’s schedule - directly altering a student’s schedule on their behalf
    • Contacting student’s parents - communicating directly with a student’s family members to address concerns or provide important information
    • Speaking with department/instructor/administrator - The advisor reaches out to a student’s department, instructor, or an administrator on behalf of them
    • Signing / filling out form for student - The advisor signs or completes a form on behalf of the student
@@
Table 13: Preliminary frequencies of the applied codes by each advisor during student-advising meetings, based on our exploratory qualitative analysis.